# Unveiling the nonlinear dynamics of a rolling axion during inflation

### v1

@@ 2.1 The model
where $\xi_{*}=\frac{\alpha\,\delta}{2}$ is the maximum value assumed by $\xi$ at the time $t=t_{*}$. We stress that these analytic results are given to leading order in slow-roll and neglectinig nackreaction from the gauge fields amplified by the motion of the axion. In particular, they assume a constant Hubble rate, and therefore the scale factor evolution $a=-\frac{1}{H\tau}$ during inflation (de Sitter geometry). In Section 4, these analytic results are compared with the full numerical solutions from the lattice.
@@ 2.2 Gauge field production
$A^{\prime\prime}_{\pm}+\left(k^{2}\pm\frac{4k\xi_{*}}{\tau\left[\left(\frac{\tau}{\tau_{*}}\right)^{\delta}+\left(\frac{\tau_{*}}{\tau}\right)^{\delta}\right]}\right)=0,$ (2.4)
@@ 2.3 Inflationary scalar perturbations
We work in the spatially flat gauge, for which (disregarding vector and tensor perturbations) the spatial components of the metric are unperturbed, $g_{ij}=a^{2}\left(\tau\right)\delta_{ij}$. The $00$ and $0i$ components of the metric are non-dynamical, and they can then be integrated out, leading to the following equations for the rescaled inflaton and axion modes $Q_{\phi}\equiv a\,\delta\phi$ and $Q_{\sigma}\equiv a\,\delta\sigma$ [31]
@@
$M^{2}_{ij}\equiv-\frac{a^{\prime\prime}}{a}\delta_{ij}+a^{2}V_{,ij}+\left(3-\frac{a^{2}\bar{\phi}^{\prime}_{k}\bar{\phi}^{\prime}_{k}}{2a^{\prime 2}}\right)\bar{\phi}^{\prime}_{i}\bar{\phi}^{\prime}_{j}+\frac{a^{3}}{a^{\prime}}(\bar{\phi}^{\prime}_{i}V_{,j}+\bar{\phi}^{\prime}_{j}V_{,i}),$ (2.8)
where the indices $i$ and $j$ run from $1$ ot $2$, and the value $1$ refers to $\phi$, while $2$ to $\sigma$ (with $\phi_{1}\equiv\phi$, $\phi_{2}\equiv\sigma$). We also note that in this relation $V$ refers to the sum of the infaton and the axion potentials, and a comma on the potential refers to differentiation wrt the corresponding field.
The last term in section 2.3 represents the direct coupling between gauge and axion modes.  [^9] As detailed in section 3, only the first equation in (2.3) is added to the system of equations solved in the lattice simulations, since the set of equations (3.1) already incorporates the coupling between gauge fields and the axion. This means that we are neglecting the backreaction of inflaton and metric perturbations on the axion-gauge field dynamics. Additionally, in the first equation of (2.3), we disregard the gravitational coupling between gauge modes and inflaton perturbations, as it is subdominant to the $\delta\phi-\delta\sigma$ coupling considered, as explained in footnote 10 of Ref. [31]. To obtain analytical predictions, one must substitute the WKB ansatz for the gauge field, derived in the previous section, into this set of equations. In appendix A, we solve the second of eqs. (2.3) and derive analytical predictions for the power spectrum of $\sigma$. The first of eqs. (2.3) is analytically solved in Ref. [31], which presents results for the power spectrum of $\phi$. In sections 4 and 5, we will compare these analytical calculations with the results from lattice simulations.
@@ 3.1 Equations of motion
where $\langle\rho\rangle$ and $\langle p\rangle$ are the background energy-density and pressure, which are obtained as lattice averages of all the energy and pressure contributions.
@@ 3.2.1 Background
The background value of the axion field is set to $(\sigma_{0},\dot{\sigma}_{0})=(0.89\,f,0.20\,f\sqrt{\bar{V}})$.  [^11] The initial values for the background inflaton, the background axion, and their time derivatives are obtained from a numerical integration of about $2$ e-folds of the exact equations for the homogeneous inflaton and axion, and for the scale factor (without including the gauge fields), that starts from slow-roll initial conditions. With this choice, there are $N\simeq 1.84$ $e$-folds left before the axion reaches its maximum velocity. The gauge field background is set to zero at the beginning of the simulation. The initial value of the scale factor $a_{0}$ is chosen so that $a=1$ when the axion reaches its maximum velocity in the simulation.
@@ 4.1 Background dynamics
We begin by presenting the evolution of background quantities, which are obtained from the simulation as averages over the $N^{3}_{\rm pts}$ lattice points. In fig. 2 we illustrate the time evolution of the background quantities $\langle\dot{\sigma}\rangle$ and $\langle\xi\rangle$, and compare them with the de Sitter approximation given by eq. 2.3, that disregards the backreaction of the gauge fields.  [^14] In comparing numerical vs. analytic results, we set $N=0$ when both the numerical and analytical values of $\dot{\sigma}$ are at their maximum. Although the discrepancy between the lattice results and the analytical approximation is small, the axion-gauge dynamics are exponentially sensitive to slight variations in $\xi$. In the following section, we will show how this difference impacts the tachyonic amplification of the gauge field.
@@ 6 Conclusions
After exploring the linear regime, we investigated the regime strongly affected by backreaction, where standard perturbation theory is not expected to provide an accurate picture. This regime is extremely relevant because many interesting observational consequences, such as a significant gravitational wave background at PTA scales, lie precisely in this regime. We found that backreaction strongly suppresses the growth of the gauge field. As a consequence, the amplitude of sourced scalar perturbations is also suppressed. Moreover, we showed that the exponential tail in the scalar distribution is suppressed by backreaction. This is due to the large number of excited gauge field modes, which tend to ‘Gaussianize’ the system due to the central limit theorem. However, we found that the suppression is less severe compared to the minimal model where the axion is the inflaton, due to the strong scale dependence in this model. Therefore, while the scale dependence does not affect the presence of an exponential tail, it influences the suppression of this tail in the strong backreaction regime. As a result, high-order non-Gaussianity can still play an important role in this system when backreaction is at play.
@@
As already mentioned, this model can contribute to the gravitational wave signal observed by PTA experiments. However, this requires working in the strong backreaction regime where analytical calculations cannot be trusted, as discussed in Ref. [59]. Although we did not calculate gravitational waves directly, we showed that backreaction significantly suppresses the growth of the gauge field in this regime, which is the primary source of the gravitational wave background. Therefore, we conclude that the parameter choices outlined in Ref. [59] are likely to result in a gravitational wave signal that is too weak to be detected by PTA experiments, due to the suppression effect observed in our simulations. Future dedicated work, employing higher-resolution lattice simulations, will be necessary to identify the viable parameter space for a detectable PTA signal. In this paper, we focused on the theoretical understanding of this system. All the phenomenological and observational prospects outlined in this Conclusion are left for future work.
@@ Acknowledgements
We are grateful to Drew Jamieson, Eiichiro Komatsu, Ippei Obata, Sébastien Renaux-Petel, Caner Ünal, and Denis Werth for insightful discussions and comments. AC acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. MP acknowledges support from Istituto Nazionale di Fisica Nucleare (INFN) through the Theoretical Astroparticle Physics (TAsP) project, and from the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - grant 20228RMX4A.

### v2

@@ 2.1 The model
where $\xi_{*}=\frac{\alpha\,\delta}{2}$ is the maximum value assumed by $\xi$ at the time $t=t_{*}$. We stress that these analytic results are given to leading order in slow-roll and neglecting backreaction from the gauge fields amplified by the motion of the axion. In particular, they assume a constant Hubble rate, and therefore the scale factor evolution $a=-\frac{1}{H\tau}$ during inflation (de Sitter geometry). In Section 4, these analytic results are compared with the full numerical solutions from the lattice.
@@ 2.2 Gauge field production
$A^{\prime\prime}_{\pm}+\left(k^{2}\pm\frac{4k\xi_{*}}{\tau\left[\left(\frac{\tau}{\tau_{*}}\right)^{\delta}+\left(\frac{\tau_{*}}{\tau}\right)^{\delta}\right]}\right)A_{\pm}=0,$ (2.4)
@@ 2.3 Inflationary scalar perturbations
We work in the spatially flat gauge, for which (disregarding vector and tensor perturbations) the line element reads
$ds^{2}=a^{2}\left(\tau\right)\left[-\left(1+2\Phi\right)d\tau^{2}+2\partial_{i}Bdx^{i}d\tau+\delta_{ij}dx^{i}dx^{j}\right]\;,$ (2.7)
where the spatial components of the metric are unperturbed, while the scalar modes $\Phi$ and $B$ are non-dynamical and can be integrated out. Namely, these modes enter in the full action $S=\int d^{4}x\sqrt{-g}\left[\frac{M_{p}^{2}}{2}R+{\cal L}\right]$ (where $R$ is the scalar curvature and ${\cal L}$ is given in eq. (2.1)) without time derivatives. Their equations of motion in momentum space are therefore constraint equations (without derivatives acting on these modes) that admit algebraic solutions for $\Phi$ and $B$ in terms of the dynamical modes $\delta\phi$ and $\delta\sigma$. We insert these solutions back into the action, which therefore becomes an action for solely the two dynamical modes. To obtain the linear equations for these modes, plus their coupling to the vector fields, we retain in this action terms that are quadratic in the dynamical modes, plus terms of ${\rm O}\left(\delta\phi\,A\right)$ and of ${\rm O}\left(\delta\sigma\,A\right)$. Finally we rescale the dynamical modes as $Q_{\phi}\equiv a\,\delta\phi$ and $Q_{\sigma}\equiv a\,\delta\sigma$, as these combinations are the canonical modes of this system (their kinetic term obtained from the procedure just outlined is $S_{\rm kin}=\frac{1}{2}\int d^{4}x\left[Q_{\phi}^{{}^{\prime}2}+Q_{\sigma}^{{}^{\prime}2}\right])$. Extremizing this action (see Section 3.1 of [31] for furhter details), gives
@@
$\tilde{M}^{2}_{ij}\equiv-\frac{a^{\prime\prime}}{a}\delta_{ij}+a^{2}V_{,ij}+\left(3-\frac{a^{2}\bar{\phi}^{\prime}_{k}\bar{\phi}^{\prime}_{k}}{2a^{\prime 2}}\right)\bar{\phi}^{\prime}_{i}\bar{\phi}^{\prime}_{j}+\frac{a^{3}}{a^{\prime}}(\bar{\phi}^{\prime}_{i}V_{,j}+\bar{\phi}^{\prime}_{j}V_{,i}),$ (2.9)
where the indices $i$ and $j$ run from $1$ ot $2$, and the value $1$ refers to $\phi$, while $2$ to $\sigma$ (with $\phi_{1}\equiv\phi$, $\phi_{2}\equiv\sigma$). We also note that in this relation $V$ refers to the sum of the inflaton and the axion potentials, and a comma on the potential refers to differentiation wrt the corresponding field.
The last term in section 2.3 represents the direct coupling between gauge and axion modes. The full structure of the “mass term” arises from the lagrangian (2.1) and from the procedure of integrating out the non-dynamical modes of the metric, which provides the gravitational interaction between the two modes (encoded in the off-diagonal $\tilde{M}^{2}_{ij}$ coefficients), as well as in their gravitational self-interactions (that contribute to the diagonal $\tilde{M}^{2}_{ij}$ coefficients).  [^9] As detailed in section 3, only the first equation in (2.3) is added to the system of equations solved in the lattice simulations, since the set of equations (3.1) already incorporates the coupling between gauge fields and the axion. This means that we are neglecting the backreaction of inflaton and metric perturbations on the axion-gauge field dynamics. Additionally, in the first equation of (2.3), we disregard the gravitational coupling between gauge modes and inflaton perturbations. This coupling, whose effect was studied in Ref. [21], is also of gravitational origin (so, at the technical level, it is equally $M_{p}$ suppressed as the one we consider here), but it occurs only at horizon crossing, while the $\delta\phi-\delta\sigma$ that we consider here continues while both fields are rolling. Comparing with the results of Ref. [21], Refs. [116] (see the final part of their Section 6.1) and [31] (see for their footnote 10) showed that the term considered here provides the dominant contribution. To obtain analytical predictions, one must substitute the WKB ansatz for the gauge field, derived in the previous section, into this set of equations. In appendix A, we solve the second of eqs. (2.3) and derive analytical predictions for the power spectrum of $\sigma$. The first of eqs. (2.3) is analytically solved in Ref. [31], which presents results for the power spectrum of $\phi$. In sections 4 and 5, we will compare these analytical calculations with the results from lattice simulations.
@@ 3.1 Equations of motion
where $\langle\rho\rangle$ and $\langle p\rangle$ are the background energy-density and pressure, which are obtained as lattice averages of all the energy and pressure contributions. Note that the gauge field does not enter the second Friedman equation explicitly (except when setting the initial conditions), as $\rho_{GF}=3p_{\rm GF}$.
@@ 3.2.1 Background
The background value of the axion field is set to $(\sigma_{0},\dot{\sigma}_{0})=(0.89\,f,0.20\,f\sqrt{\bar{V}})$.  [^11] The initial values for the background inflaton, the background axion, and their time derivatives are obtained from a numerical integration of about $2$ e-folds of the exact equations for the homogeneous inflaton and axion, and for the scale factor (without including the gauge fields), that starts from slow-roll initial conditions. The initial condition for the axion is tuned so that it reaches its maximum velocity at $N\simeq 1.84$ $e$-folds after the start of the simulation. The gauge field background is set to zero at the beginning of the simulation. The initial value of the scale factor $a_{0}$ is chosen so that $a=1$ when the axion reaches its maximum velocity in the simulation.
@@ 4.1 Background dynamics
We begin by presenting the evolution of background quantities, which are obtained from the simulation as averages over the $N^{3}_{\rm pts}$ lattice points. In fig. 2 we illustrate the time evolution of the background quantities $\langle\dot{\sigma}\rangle$ and $\langle\xi\rangle$, and compare them with the de Sitter approximation given by eq. 2.3, that disregards both backreaction of the gauge fields and corrections beyond the slow-roll approximation.  [^14] In comparing numerical vs. analytic results, we set $N=0$ when both the numerical and analytical values of $\dot{\sigma}$ are at their maximum. Although the discrepancy between the lattice results and the analytical approximation is small, the axion-gauge dynamics are exponentially sensitive to slight variations in $\xi$. In the following section, we will show how this difference impacts the tachyonic amplification of the gauge field.
@@ 6 Conclusions
After exploring the linear regime, we investigated the regime strongly affected by backreaction, where standard perturbation theory is not expected to provide an accurate picture. This regime is extremely relevant because many interesting observational consequences, such as a significant gravitational wave background at PTA scales, lie precisely in this regime. We found that backreaction strongly suppresses the growth of the gauge field. As a consequence, the amplitude of sourced scalar perturbations is also suppressed. Still, we see from Figure 16 that the sourced perturbations are greater than the vacuum ones in the regime of strong backreaction. Although the case we studied has significant differences with that of an axion inflation, it is hard to imagine that a prolonged motion of the axion throughout inflation, and a direct coupling to the gauge fields of the field responsible for the density perturbations will result in a smaller sourced signal than the one emerged from our study. Therefore our results confirm that a regime of strong backreaction in axion inflation [78] can be at play only well after the CMB modes have been produced.
Concerning the statistics of the perturbations, we showed that the exponential tail in the scalar distribution is suppressed by backreaction. This is due to the large number of excited gauge field modes, which tend to ‘Gaussianize’ the system due to the central limit theorem. However, we found that the suppression is less severe compared to the minimal model where the axion is the inflaton, due to the strong scale dependence in this model. Therefore, while the scale dependence does not affect the presence of an exponential tail, it influences the suppression of this tail in the strong backreaction regime. As a result, high-order non-Gaussianity can still play an important role in this system when backreaction is at play.
@@
As already mentioned, this model can contribute to the gravitational wave signal observed by PTA experiments. However, this requires working in the strong backreaction regime where analytical calculations cannot be trusted, as discussed in Ref. [59]. Although we did not calculate gravitational waves directly, we showed that backreaction significantly suppresses the growth of the gauge field in this regime, which is the primary source of the gravitational wave background. Therefore, we conclude that the parameter choices outlined in Ref. [59] are likely to result in a gravitational wave signal that is too weak to be detected by PTA experiments, due to the suppression effect observed in our simulations. Future dedicated work, employing higher-resolution lattice simulations, will be necessary to identify the viable parameter space for a detectable PTA signal. In this paper, we focused on the theoretical understanding of this system. All the phenomenological and observational prospects outlined in these Conclusions are left for future work.
@@ Acknowledgements
We are grateful to Drew Jamieson, Eiichiro Komatsu, Ippei Obata, Sébastien Renaux-Petel, Lorenzo Sorbo, Caner Ünal, and Denis Werth for insightful discussions and comments. AC acknowledges funding support from the Initiative Physique des Infinis (IPI), a research training program of the Idex SUPER at Sorbonne Université. MP acknowledges support from Istituto Nazionale di Fisica Nucleare (INFN) through the Theoretical Astroparticle Physics (TAsP) project, and from the MIUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - grant 20228RMX4A.